OpenStudy (anonymous):

What is the value of 7P3? A. 35 B. 56 C. 210 D. 840 WILL GIVE MEDAL!!!!!!!! PLEASE HELP

ganeshie8 (ganeshie8):

7P3 = 7*6*5 = 7*30 = 210

OpenStudy (anonymous):

thanks @ganeshie8